# Nonequilibrium steady states in multi-bath quantum collision models

### v2

@@ I Introduction
A framework for understanding how a quantum system evolves when in contact with an environment is critical; both in terms of the practical development of quantum technologies and in addressing long-standing open questions regarding the dynamics of complex systems. Regarding the former noise characterisation in prototype quantum devices is crucial to overcome scalability issues ManiscalcoScalability, while the latter touches on major conceptual gaps in our understanding, e.g. determining the quantum-to-classical transition ZurekRMP or how complex systems thermalise Deutsch2018. It is not surprising then that a range of tools have been developed to simulate open quantum systems PReB; Ferracin2024; LandiRMP; Breuer2002; GooldPRX. Within this plethora of techniques, collision models CICCARELLO20221; Campbell2021EPL; cusumano2022quantum; Rau1963; ScaraniPRL2002; Homogenisation or repeated interaction schemes StrasbergPRX, stand as a particularly versatile approach having been used to reconcile apparent thermodynamic inconsistencies present in other approaches DeChiaraNJP, explore how classicality emerges from an otherwise quantum dynamics CampbellPRA2019; Ryan2022, explore synchronization in quantum systems karpat2019pra, and efficiently cool quantum systems HuberCooling; KochCooling.
@@
The simplest formulation of a collision model involves imagining the environment as an infinite collection of identically prepared constituents, here after referred to as “auxiliary units” and labelled $A$, with which the system, $S$, sequentially interacts. Depending on the precise form of the $S$-AA interaction, and the initial states of the auxiliary units, this model accurately captures a range of Markovian dynamics. In particular, for so-called “thermal operations” it leads to thermalisation ScaraniPRL2002, while the model also more generally leads to homogenisation Homogenisation. However, varying the $S$-AA interaction can lead to a range of non-equilibrium steady states (NESS), whose characterisation has recently become the focus of several works Guarnieri2020PLA; Heineken_NESS_Entanglement; Tian_Heat_Transfer_2021; Segal2025QST; Baris1; Baris2; Hammam2024quantum; cusumano2024structured; hammam2021optimizing; corr2025continuous. While assuming that all the auxiliary units are identically prepared is well justified, particularly in order to simplify simulation, this is clearly not a very realistic model. For instance, a more precise microscopic description of a thermal environment would assume that the system interacts with auxiliary units whose initial state is drawn in accordance with a distribution of energy that correspond to a given temperature KochCooling. Several works have explored relaxing the assumption of identical auxiliary units and deterministic collisional events, e.g. by leveraging scattering techniques Scattering1; Scattering2; Scattering3. The flexibility of collision models is clearly one of their inherent strengths that allows us to explore a range of other open system dynamics Lorenzo2015heat; cusumano2017interferometric; cusumano2018interferometric; pedram2022environment; Erbanni2023simulating; Garg2025simulating. From the preceding discussion it is clear that collision models provide a useful approach to simulate how a system reaches the same, possibly thermal or equilibrium, steady state from different microscopic models.
@@
This serves as the starting point for our analysis. We model a simple thermalisation process via collision models for two settings: the ‘typical’ situation where the auxiliary units are assumed to be all identical thermal states and a second setting where the system interacts simultaneously with two collisional baths with different initial energies, and these energies dictate the strength of the $S$-AA interaction. While these approaches are known to recover the expected Markovian dynamics as described by the standard GKSL master equation CICCARELLO20221; cusumano2022quantum (see the Appendix), we show that they correspond to physically distinct processes, with the latter setting clearly driving the system to a NESS. Through a simple modification of the basic collision model, we examine how the introduction of non-Markovianity to the dynamics significantly changes the steady state properties observed, with the establishment of strong system bath correlations playing a particularly important role. We explore the impact of introducing projective measurements on the auxiliary units using a two-point measurement scheme to recover the average heat current in the presence of non-Markovianity and briefly comment on the impact that alternative system-environment interaction terms can have on the steady-state properties.
@@ II.1 Single-bath thermalizing model (setting I)
Collision models are a flexible framework for studying open system dynamics CICCARELLO20221; Campbell2021EPL; indeed they can be used to simulate any time-continuous open system dynamics described by a Markovian quantum master equation MEsfromCMs; Cattaneo2021collision. A simple setting of wide physical relevance is a two-level system with Hamiltonian, $H_{S}=\tfrac{\omega_{S}}{2}\sigma_{z}$, thermalising with a bosonic environment described by the master equation Current_fluctuations,
@@
where $g=(2\bar{N}+1)^{-1}$ 2-level_thermalization.
The dynamics described by Eq. (4) can be stroboscopically simulated using a collision model Rau1963; ScaraniPRL2002; CICCARELLO20221; cusumano2022quantum, where the environment is viewed as an infinite collection of auxiliary units, each modeled as a qubit, identically prepared in a thermal state, i.e. $\eta_{n}=\rho^{\text{th}}$ given by Eq. (5), with the free evolution of each unit given by $H_{A_{n}}=\tfrac{\omega_{A_{n}}}{2}\sigma^{z}$. For simplicity, in the rest of this work we shall always assume the system and auxiliary units are on resonance, i.e. $\omega_{S}\!=\!\omega_{A_{n}}\!\equiv\!\omega$. This setting is illustrated in Fig. 1(b). In order to match the dynamics given by Eq. (4), we fix the interaction Hamiltonian between the $S$ and $A_{n}$ to be,
@@
with $J=\sqrt{{\frac{\Gamma(2\bar{N}+1)}{\Delta t}}}$. Scaling the interaction strength by $1/\sqrt{\Delta t}$ is necessary to ensure that Eq. (4) is recovered in the limit $\Delta t\to 0$. Throughout this work we will refer to this as ‘setting I’. Henceforth, we shall move to the interaction picture and neglect the free evolution of the system and auxiliary units. This can be done due to the fact that sum of the free Hamiltonians commute with the interaction Hamiltonian, i.e. $\left[H_{S}+H_{A_{n}},H_{I}\right]=0$  Composite_CM. This fact also means there is no switching work required to stoboscopically turn on and off the system-bath interaction DeChiaraNJP; Guarnieri2020PLA.
@@
This setup leads to thermalization, wherein the steady state of $S$ converges to the initial auxiliary state, i.e. $\rho_{\infty}\!=\!\eta_{i}$, for any choice of initial state for the auxiliary units $A_{n}$ Homogenisation. When $S$ sequentially interacts with $A_{n}$ via Eq. (6), its dynamics are described by a contractive completely positive map, guaranteeing convergence toward $\eta$ as the unique fixed point. This process is thermodynamically consistent and captures key features of thermalization. Importantly, in the limit of infinitely many collisions, the system not only reaches equilibrium but also asymptotically factorizes from the environment, resulting in a product state without residual system-environment correlations GiovannettiPRA2018.
While here we will focus on system-environment interactions of the form Eq. (6), in the Appendix we consider an alternative given by the isotropic Heisenberg Hamiltonian. Although this interaction also leads to homogenization in setting I, the collision model no longer reduces to the same thermalisation dynamics governed by the GKSL equation, Eq. (4cusumano2024structured. Indeed, while many of the main features discussed below are similarly exhibited for the Heisenberg interaction, we highlight a number of interesting differences that emerge when alternative methods for simulating the open dynamics are considered in both the Markovian and non-Markovian regimes.
@@ II.3 Steady States in the Markovian Limit
Maintaining the system in a NESS requires heat currents to be flowing. Because the unitary is strictly energy preserving, there is no work associated with switching on and off the collisional unitaries DeChiaraNJP; Guarnieri2020PLA. Hence all energy changes can be considered as heat exchanges. Focusing on the zero temperature bath in setting II, within the collision model setting we can readily calculate this via
@@
where $\tilde{\rho}_{A_{n}^{(0)}}$ is the state of the $A_{n}^{(0)}$ auxiliary unit after the system-environment collision has taken place. A similar expression holds for the $A_{n}^{(1)}$’s and, in fact, we find is exactly equal and opposite to Eq. (9), which is a direct consequence of strict energy conservation. Fig. 2 shows the steady state heat flux into the environment, i.e. $\frac{\Delta\mathcal{Q}}{\Delta t}\!\equiv\!\frac{\Delta\mathcal{Q}^{(0)}(n)}{\Delta t}$ for $n\!\rightarrow\!\infty$. We see that it is non-zero for all values of $\Delta t$. Smaller values of $\beta$ require larger heat fluxes to maintain the NESS. This is in line with expectation since in this regime the system is being driven to a state with a high effective temperature. It is therefore far from equilibrium from both of the respective collisional baths leading to substantial energy exchanges. Larger $\beta$ leads to a reduced heat flux since this regime tends to the zero temperature single bath situation captured by setting I. It is worth noting that the heat flux vanishes for all choices of $\beta$ in setting I Guarnieri2020PLA.
@@ III Non-Markovian Quantum Collision Models
When using collision models to simulate Markovian dynamics, we make three assumptions regarding the constituents of the bath: (i) they do not interact with each other; (ii) they are initially uncorrelated; (iii) each constituent interacts with $S$ only once. Relaxing these assumptions allows us to extend collision models to simulate non-Markovian dynamics ruariMauro; csenyacsa2022entropy; BarisPRA2017; Pezzutto_Paternostro_2016; Zhang2019; Pleasance_2025.
@@
The impact of intra-environment interactions is to correlate $S$ with $A_{n+1}$ before they have directly interacted with each other. This leads to a non-Markovian dynamics NmReview with the parameter $\delta$ controlling the strength of the non-Markovianity ruariMauro. For $\delta\!=\!0$ the dynamics reverts to the Markovian case as discussed in Sec. II.3, while $\delta\!=\!\frac{\pi}{2}$ leads to an infinitely non-Markovian dynamics ruariMauro since, for this interaction strength, Eq. (10) becomes a full SWAP and therefore the system is effectively interacting with the same qubit in the environment at every timestep and there is no loss of information.
A useful way to understand the introduction of the intra-environment interactions is through the notion of a memory-depth Bassano. For the Markovian collision models in Sec. II.3, each environmental constituent is involved in only a single collision and, thus, can be traced out after this interaction has taken place since it plays no role in the ensuing dynamics. Clearly, this is not the case for the non-Markovian setting as each $A_{n}$ is involved in two collisions, one with the system and a second with the incoming bath qubit, $A_{n+1}$. Introducing an additional full SWAP operation applied to $A_{n}$ and $A_{n+1}$ in Eqs. (12) allows to recover an effective Markovian dynamics Bassano; AA_to_Composite. This is because the additional SWAP operation means that, in effect, the system interacts with the same bath unit, now considered a memory, during every collision. While this may seem like an unnecessary complication, it means that at the level of the system plus memory compound the dynamics is Markovian Bassano; Composite_CM; AA_to_Composite. This Markovian embedding of non-Markovian dynamics allows to readily determine the system-memory steady state, from which the steady state of the system directly follows by tracing over the memory’s degrees of freedom.
Our interest is in characterising the steady states arising from the different microscopic models captured by Fig. 1. The ability to introduce non-Markovianity via intra-environment interactions therefore adds an interesting additional dimension to our study, complementing recent works that focused on the Markovian limit Guarnieri2020PLA; Segal2025QST. To this end, we aim to assess the impact, if any, that degree of non-Markovianity plays in the resulting steady state properties. Several metrics have been proposed to quantify non-Markovianity NmReview; NMmeasures. We shall focus on the BLP measure BLPmeasure; BLPmeasure2, which takes an information theoretic approach based on the behaviour of the trace distance between two states,
@@
where $\Omega_{+}=\bigcup_{i}(a_{i},b_{i})$ is the union of all the time intervals in the dynamics in which the trace distance between $\rho$ and $\pi$ increases, i.e. $\partial_{t}D(\rho(t),\pi(t))>0$. Here, $\rho(t)$ and $\pi(t)$ start from initially distinguishable (orthogonal) states. Equation (14) can be interpreted as summing over all the regions where the trace distance between a pair of system states increases during the dynamics, indicating an increase in their distinguishability and corresponding to an information backflow from the environment. The maximization is taken over all possible pairs of initial states. However, it has been shown that for qubit states, as is the focus of this work, the states that maximise Eq. (14) are always pure, orthogonal states OptimalNM_states. For a collision model which simulates the open system dynamics via discrete time steps, we shall use a discretized version of Eq. (14) such that the non-Markovianity is quantified as ruariMauro; BLPmeasure2,

### v3

@@ I Introduction
A framework for understanding how a quantum system evolves when in contact with an environment is critical; both in terms of the practical development of quantum technologies and in addressing long-standing open questions regarding the dynamics of complex systems. Regarding the former noise characterisation in prototype quantum devices is crucial to overcome scalability issues Filippov et al. (2024), while the latter touches on major conceptual gaps in our understanding, e.g. determining the quantum-to-classical transition Zurek (2003) or how complex systems thermalise Deutsch (2018). It is not surprising then that a range of tools have been developed to simulate open quantum systems Purkayastha et al. (2021); Ferracin et al. (2024); Landi et al. (2022); Breuer and Petruccione (2002); Brenes et al. (2020). Within this plethora of techniques, collision models Ciccarello et al. (2022); Campbell and Vacchini (2021); Cusumano (2022); Rau (1963); Scarani et al. (2002); Ziman et al. (2002) or repeated interaction schemes Strasberg et al. (2017), stand as a particularly versatile approach having been used to reconcile apparent thermodynamic inconsistencies present in other approaches De Chiara et al. (2018), explore how classicality emerges from an otherwise quantum dynamics Campbell et al. (2019); Ryan et al. (2022), explore synchronization in quantum systems Karpat et al. (2019), and efficiently cool quantum systems Taranto et al. (2020); Langbehn et al. (2025).
@@
The simplest formulation of a collision model involves imagining the environment as an infinite collection of identically prepared constituents, here after referred to as “auxiliary units” and labelled $A$, with which the system, $S$, sequentially interacts. Depending on the precise form of the $S$-AA interaction, and the initial states of the auxiliary units, this model accurately captures a range of Markovian dynamics. In particular, for so-called “thermal operations” it leads to thermalisation Scarani et al. (2002), while the model also more generally leads to homogenisation Ziman et al. (2002). However, varying the $S$-AA interaction can lead to a range of non-equilibrium steady states (NESS), whose characterisation has recently become the focus of several works Guarnieri et al. (2020); Heineken et al. (2021); Tian et al. (2021); Prositto et al. (2025a); Palafox et al. (2022); Santiago-García et al. (2025); Hammam et al. (2024); Cusumano and De Chiara (2024); Hammam et al. (2021); Corr et al. (2025). While assuming that all the auxiliary units are identically prepared is well justified, particularly in order to simplify simulation, this is clearly not a very realistic model. For instance, a more precise microscopic description of a thermal environment would assume that the system interacts with auxiliary units whose initial state is drawn in accordance with a distribution of energy that correspond to a given temperature Langbehn et al. (2025). Several works have explored relaxing the assumption of identical auxiliary units and deterministic collisional events, e.g. by leveraging scattering techniques Ehrich et al. (2020); Jacob et al. (2021); Tabanera-Bravo et al. (2023). The flexibility of collision models is clearly one of their inherent strengths that allows us to explore a range of other open system dynamics Lorenzo et al. (2015); Cusumano et al. (2017, 2018b); Pedram et al. (2022); Erbanni et al. (2023); Garg et al. (2025). From the preceding discussion it is clear that collision models provide a useful approach to simulate how a system reaches the same, possibly thermal or equilibrium, steady state from different microscopic models.
@@
This serves as the starting point for our analysis. Our aim is to critically explore the differences between collision models involving a single thermal bath and those featuring multiple pure-state baths, particularly in the non-Markovian regime. We start by considering a simple thermalisation process via collision models for two settings: the ‘typical’ situation where the auxiliary units are assumed to be all identical thermal states and a second setting where the system interacts simultaneously with two collisional baths with different initial energies, and these energies dictate the strength of the $S$-AA interaction. While these approaches are known to recover the expected Markovian dynamics as described by the standard GKSL master equation Ciccarello et al. (2022); Cusumano (2022) (see the Appendix), we show that they correspond to physically distinct processes, with the latter setting clearly driving the system to a NESS. Through a simple modification of the basic collision model, we examine how the introduction of non-Markovianity to the dynamics significantly changes the steady state properties observed, with the establishment of strong system bath correlations playing a particularly important role. We explore the impact of introducing projective measurements on the auxiliary units using a two-point measurement scheme to recover the average heat current in the presence of non-Markovianity and briefly comment on the impact that alternative system-environment interaction terms can have on the steady-state properties.
@@ II.1 Single-bath thermalizing model (setting I)
Collision models are a flexible framework for studying open system dynamics Ciccarello et al. (2022); Campbell and Vacchini (2021); indeed they can be used to simulate any time-continuous open system dynamics described by a Markovian quantum master equation Ziman et al. (2005); Cattaneo et al. (2021). A simple setting of wide physical relevance is a two-level system with Hamiltonian, $H_{S}=\tfrac{\omega_{S}}{2}\sigma_{z}$, thermalising with a bosonic environment described by the master equation Landi et al. (2024),
@@
where $g=(2\bar{N}+1)^{-1}$ Cherian et al. (2019).
The dynamics described by Eq. (4) can be stroboscopically simulated using a collision model Rau (1963); Scarani et al. (2002); Ciccarello et al. (2022); Cusumano (2022), where the environment is viewed as an infinite collection of auxiliary units, each modeled as a qubit, identically prepared in a thermal state, i.e. $\eta_{n}=\rho^{\text{th}}$ given by Eq. (5), with the free evolution of each unit given by $H_{A_{n}}=\tfrac{\omega_{A_{n}}}{2}\sigma^{z}$. For simplicity, in the rest of this work we shall always assume the system and auxiliary units are on resonance, i.e. $\omega_{S}\!=\!\omega_{A_{n}}\!\equiv\!\omega$. This setting is illustrated in Fig. 1(b). In order to match the dynamics given by Eq. (4), we fix the interaction Hamiltonian between the $S$ and $A_{n}$ to be,
@@
with $J=\sqrt{{\frac{\Gamma(2\bar{N}+1)}{\Delta t}}}$. Scaling the interaction strength by $1/\sqrt{\Delta t}$ is necessary to ensure that Eq. (4) is recovered in the limit $\Delta t\to 0$. Throughout this work we will refer to this as ‘setting I’. Henceforth, we shall move to the interaction picture and neglect the free evolution of the system and auxiliary units. This can be done due to the fact that sum of the free Hamiltonians commute with the interaction Hamiltonian, i.e. $\left[H_{S}+H_{A_{n}},H_{I}\right]=0$  Lorenzo et al. (2017). This fact also means there is no switching work required to stoboscopically turn on and off the system-bath interaction De Chiara et al. (2018); Guarnieri et al. (2020).
@@
This setup leads to thermalization, wherein the steady state of $S$ converges to the initial auxiliary state, i.e. $\rho_{\infty}\!=\!\eta_{i}$, for any choice of initial state for the auxiliary units $A_{n}$ Ziman et al. (2002). When $S$ sequentially interacts with $A_{n}$ via Eq. (6), its dynamics are described by a contractive completely positive map, guaranteeing convergence toward $\eta$ as the unique fixed point. This process is thermodynamically consistent and captures key features of thermalization. Importantly, in the limit of infinitely many collisions, the system not only reaches equilibrium but also asymptotically factorizes from the environment, resulting in a product state without residual system-environment correlations Cusumano et al. (2018a).
While here we will focus on system-environment interactions of the form Eq. (6), in the Appendix we consider an alternative given by the isotropic Heisenberg Hamiltonian. Although this interaction also leads to homogenization in setting I, the collision model no longer reduces to the same thermalisation dynamics governed by the GKSL equation, Eq. (4Cusumano and De Chiara (2024). Indeed, while many of the main features discussed below are similarly exhibited for the Heisenberg interaction, we highlight a number of interesting differences that emerge when alternative methods for simulating the open dynamics are considered in both the Markovian and non-Markovian regimes.
@@ II.3 Steady States in the Markovian Limit
Maintaining the system in a NESS requires heat currents to be flowing. Because the unitary is strictly energy preserving, there is no work associated with switching on and off the collisional unitaries De Chiara et al. (2018); Guarnieri et al. (2020). Hence all energy changes can be considered as heat exchanges. Focusing on the zero temperature bath in setting II, within the collision model setting we can readily calculate this via
@@
where $\tilde{\rho}_{A_{n}^{(0)}}$ is the state of the $A_{n}^{(0)}$ auxiliary unit after the system-environment collision has taken place. A similar expression holds for the $A_{n}^{(1)}$’s and, in fact, we find is exactly equal and opposite to Eq. (9), which is a direct consequence of strict energy conservation. Fig. 2 shows the steady state heat flux into the environment, i.e. $\frac{\Delta\mathcal{Q}}{\Delta t}\!\equiv\!\frac{\Delta\mathcal{Q}^{(0)}(n)}{\Delta t}$ for $n\!\rightarrow\!\infty$. We see that it is non-zero for all values of $\Delta t$. Smaller values of $\beta$ require larger heat fluxes to maintain the NESS. This is in line with expectation since in this regime the system is being driven to a state with a high effective temperature. It is therefore far from equilibrium from both of the respective collisional baths leading to substantial energy exchanges. Larger $\beta$ leads to a reduced heat flux since this regime tends to the zero temperature single bath situation captured by setting I. It is worth noting that the heat flux vanishes for all choices of $\beta$ in setting I Guarnieri et al. (2020).
@@ III Non-Markovian Quantum Collision Models
When using collision models to simulate Markovian dynamics, we make three assumptions regarding the constituents of the bath: (i) they do not interact with each other; (ii) they are initially uncorrelated; (iii) each constituent interacts with $S$ only once. Relaxing these assumptions allows us to extend collision models to simulate non-Markovian dynamics McCloskey and Paternostro (2014); Şenyaşa et al. (2022); Çakmak et al. (2017); Pezzutto et al. (2016); Zhang et al. (2019); Pleasance et al. (2025).
@@
The impact of intra-environment interactions is to correlate $S$ with $A_{n+1}$ before they have directly interacted with each other. This leads to a non-Markovian dynamics Breuer et al. (2016) with the parameter $\delta$ controlling the strength of the non-Markovianity McCloskey and Paternostro (2014). For $\delta\!=\!0$ the dynamics reverts to the Markovian case as discussed in Sec. II.3, while $\delta\!=\!\frac{\pi}{2}$ leads to an infinitely non-Markovian dynamics McCloskey and Paternostro (2014) since, for this interaction strength, Eq. (10) becomes a full SWAP and therefore the system is effectively interacting with the same qubit in the environment at every timestep and there is no loss of information.
A useful way to understand the introduction of the intra-environment interactions is through the notion of a memory-depth Campbell et al. (2018). For the Markovian collision models in Sec. II.3, each environmental constituent is involved in only a single collision and, thus, can be traced out after this interaction has taken place since it plays no role in the ensuing dynamics. Clearly, this is not the case for the non-Markovian setting as each $A_{n}$ is involved in two collisions, one with the system and a second with the incoming bath qubit, $A_{n+1}$. Introducing an additional full SWAP operation applied to $A_{n}$ and $A_{n+1}$ in Eqs. (12) allows to recover an effective Markovian dynamics Campbell et al. (2018); Kretschmer et al. (2016a). This is because the additional SWAP operation means that, in effect, the system interacts with the same bath unit, now considered a memory, during every collision. While this may seem like an unnecessary complication, it means that at the level of the system plus memory compound the dynamics is Markovian Campbell et al. (2018); Lorenzo et al. (2017); Kretschmer et al. (2016a). This Markovian embedding of non-Markovian dynamics allows to readily determine the system-memory steady state, from which the steady state of the system directly follows by tracing over the memory’s degrees of freedom.
Our interest is in characterising the steady states arising from the different microscopic models captured by Fig. 1. The ability to introduce non-Markovianity via intra-environment interactions therefore adds an interesting additional dimension to our study, complementing recent works that focused on the Markovian limit Guarnieri et al. (2020); Prositto et al. (2025a). To this end, we aim to assess the impact, if any, that degree of non-Markovianity plays in the resulting steady state properties. Several metrics have been proposed to quantify non-Markovianity Breuer et al. (2016); Addis et al. (2014). We shall focus on the BLP measure Breuer et al. (2009); Laine et al. (2010), which takes an information theoretic approach based on the behaviour of the trace distance between two states,
@@
where $\Omega_{+}=\bigcup_{i}(a_{i},b_{i})$ is the union of all the time intervals in the dynamics in which the trace distance between $\rho$ and $\pi$ increases, i.e. $\partial_{t}D(\rho(t),\pi(t))>0$. Here, $\rho(t)$ and $\pi(t)$ start from initially distinguishable (orthogonal) states. Equation (14) can be interpreted as summing over all the regions where the trace distance between a pair of system states increases during the dynamics, indicating an increase in their distinguishability and corresponding to an information backflow from the environment. The maximization is taken over all possible pairs of initial states. However, it has been shown that for qubit states, as is the focus of this work, the states that maximise Eq. (14) are always pure, orthogonal states Wißmann et al. (2012). For a collision model which simulates the open system dynamics via discrete time steps, we shall use a discretized version of Eq. (14) such that the non-Markovianity is quantified as McCloskey and Paternostro (2014); Laine et al. (2010),